I really need help with this one question,
it asks, ''What is the solution of the proportion,
x-1/2 = x-4/5?
The letters are,
A. -2.33; B. 1; C. -1.86; D. -1
Please help? It's only this one.

Answer by jim_thompson5910(35256)

(x-1)/2 = (x-4)/5

5(x-1) = 2(x-4) ... cross multiply

5x-5 = 2x-8

5x = 2x-8+5

5x-2x = -8+5

3x = -3

x = -3/3

x = -1